OpenStudy (hahawowoh):

Will fan and medal! Simplify (5x2 + 3x + 4) + (2x2 − 6x + 3).

OpenStudy (texaschic101):

since there is a + in between, the parenthesis are not needed. 5x^2 + 3x + 4 + 2x^2 - 6x + 3 now we will combine like terms 5x^2 + 2x^2 = 7x^2 3x - 6x = -3x 4 + 3 = 7 so that leaves us with : 7x^2 - 3x + 7 questions ?

OpenStudy (hahawowoh):

Mo, thank you so much!